# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Appendix A Appendix: PXP Model
$\begin{split}\hat{H}_{\mathrm{FLH}}=&\hat{H}_{+}+\hat{H}_{-}+\hat{H}_{\mathrm{v}}+\hat{H}_{\mathrm{x}},\\ \hat{H}_{\pm}=&J(1-\eta)\sum^{N}_{j=1}\left[\left(\hat{\sigma}^{+}_{j,{\pm}}\hat{\sigma}^{-}_{j+1,\pm}+\mathrm{h.c.}\right)+\hat{\sigma}^{z}_{j,{\pm}}\hat{\sigma}^{z}_{j+1,\pm}\right],\\ \hat{H}_{\mathrm{v}}=&\frac{1}{2}J\sum^{N}_{j=1}\left[\left(\hat{\sigma}^{+}_{j,-}\hat{\sigma}^{-}_{j,+}+\mathrm{h.c.}\right)+\hat{\sigma}^{z}_{j,-}\hat{\sigma}^{z}_{j,+}\right],\\ \hat{H}_{\mathrm{x}}=&J(1+\eta)\sum^{N}_{j=1}\left[\left(\hat{\sigma}^{+}_{j,\pm}\hat{\sigma}^{-}_{j+1,\mp}+\mathrm{h.c.}\right)+\hat{\sigma}^{z}_{j,\pm}\hat{\sigma}^{z}_{j+1,\mp}\right].\end{split}$ (S34)
@@
In Fig. S5, we see that when the local HSF is not coupled with the thermalization part, the dephasing system remains in a state of LPTS breaking, while conversely, a clear process of LPTS breaking occurs. It should be noted that the red dots in Fig. S5(a)-(b) represent a few typical local HSF states and do not include all local HSF states.
Appendix S8 F. Detail of the imbalance dynamics
The LPTS breaking can also be manifested through the dynamical behavior of the density imbalance, $I(t)=(1/L)\sum_{j}\langle\!\langle\rho(0)|\hat{\sigma}^{z}_{j}\otimes\hat{\sigma}^{z}_{j}|\rho(0)\rangle\!\rangle\langle\!\langle\rho(t)|\hat{\sigma}^{z}_{j}\otimes\hat{\sigma}^{z}_{j}|\rho(t)\rangle\!\rangle$, for some initial state $|\rho(0)\rangle\!\rangle$. The imbalance dynamics in dissipation-free scenarios exhibit the form of a damped oscillation for special (scarred) initial states, i.e., $|\Pi\rangle\otimes|\Pi\rangle$ state in the SCL model and $|\mathbb{Z}_{2}\rangle\otimes\mathbb{Z}_{2}\rangle$ state in the PXP model. These QMBS oscillations can be fitted by using $\bar{I}(t)=\exp(-\beta t)\cos(\omega t)$ over one time periods in Figs. S6(a)-(b).
Figure S6: (a)-(b): The time evolution of density imbalance $I(t)$ for SCL and PXP models for different pure dephasing rates $\gamma$. The gray dashed-dot represents the fitting function $\bar{I}(t)$. The coefficients $(\beta,\omega)$ of the imblance fitting function $\bar{I}(t)$ are shown in panel. The system size is $L=14$.

### v3

@@ Abstract
$\begin{split}\hat{H}_{\mathrm{FLH}}=&\hat{H}_{+}+\hat{H}_{-}+\hat{H}_{\mathrm{v}}+\hat{H}_{\mathrm{x}},\quad\quad\quad\quad\mathrm{with}\quad\quad\hat{H}_{\pm}=J(1-\eta)\sum^{N}_{j=1}\left[\left(\hat{\sigma}^{+}_{j,{\pm}}\hat{\sigma}^{-}_{j+1,\pm}+\mathrm{h.c.}\right)+\hat{\sigma}^{z}_{j,{\pm}}\hat{\sigma}^{z}_{j+1,\pm}\right],\\ \hat{H}_{\mathrm{v}}=&\frac{1}{2}J\sum^{N}_{j=1}\left[\left(\hat{\sigma}^{+}_{j,-}\hat{\sigma}^{-}_{j,+}+\mathrm{h.c.}\right)+\hat{\sigma}^{z}_{j,-}\hat{\sigma}^{z}_{j,+}\right],\quad\hat{H}_{\mathrm{x}}=J(1+\eta)\sum^{N}_{j=1}\left[\left(\hat{\sigma}^{+}_{j,\pm}\hat{\sigma}^{-}_{j+1,\mp}+\mathrm{h.c.}\right)+\hat{\sigma}^{z}_{j,\pm}\hat{\sigma}^{z}_{j+1,\mp}\right].\end{split}$ (S61)
@@
In Fig. S6, we see that when the local HSF is not coupled with the thermalization part, the dephasing system remains in a state of Liouvillian spectrum transition, while conversely, a clear process of Liouvillian spectrum transition occurs. It should be noted that the red dots in Fig. S6(a)-(b) represent a few typical local HSF states and do not include all local HSF states.
S6 F. Details of fitting the imbalance dynamics
Liouvillian spectrum transition can also be manifested through the dynamical behavior of the density imbalance $I(t)$ for the special initial state $|\rho(0)\rangle\!\rangle$, i.e., $|\Pi\rangle\otimes|\Pi\rangle$ state in the CL model. These QMBS oscillations can be fitted by using $\bar{I}(t)=\exp(-\beta t)\cos(\omega t)$ over one or several time periods, and Fig. S7 demonstrates the robustness of such fits. Through our numerical analysis, we found that compared to fitting several revivals, fitting one period results in a slight shift for the decay coefficient $\beta$, but the Liouvillian spectrum transition point remains unchanged. Since it is numerically costly to obtain data over many periods in large system sizes, in the main text we have restricted our analysis to the first period.
Figure S7: The coefficients $(\beta,\omega)$ of the imbalance fitting function $\bar{I}(t)$ as a function of dephasing $\gamma L$ for the CL model for different fitting intervals $t\in[0,t_{\rm{max}}]$ with $t_{\rm{max}}=3,6,12$. All data is for $L=14$, $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!\in\![0,0.2]$ drawn from a uniform distribution.